OpenStudy (ny,ny):

Find all open intervals on which (x^2)/(x^2+4) is decreasing.

OpenStudy (steve816):

Step 1: Take the derivative Step 2: Solve for x when the derivative = 0 Step 3: Interval test to see when the derivative is negative or the derivative is positive. When the derivative is negative, that means that the function is decreasing.

OpenStudy (steve816):

Let me know if you get stuck.

OpenStudy (solomonzelman):

Steve, I think it would be better to just set f'(x)<0, for step 2.

OpenStudy (solomonzelman):

(of course we all agree a derivative is necessary) ... NY NY, have you found the derivative?

OpenStudy (ny,ny):

I got x=0 And to do the interval test, do i plug in a number less than and greater than 0?

OpenStudy (solomonzelman):

What was the derivative of the function?

OpenStudy (ny,ny):

I got (8x)/((x+4)^2)

OpenStudy (solomonzelman):

You mean (x^2+4)^2 ?

OpenStudy (steve816):

@SolomonZelman I guess you could set f'(x) < 0, but to find the values, I think you are essentially doing the same thing by interval tests.

OpenStudy (ny,ny):

Oh yes, i wrote it wrong.

OpenStudy (solomonzelman):

Yes ... I'm just giving a suggestion, because as soon as you set f'(x)<0, it becomes clear.

OpenStudy (solomonzelman):

The function is decreasing when the slope (or, the derivative) is negative, right?

OpenStudy (solomonzelman):

For this reason, to see where the function is decreasing, we set \(f'(x)<0\). Or, in your case, \(\color{black}{\displaystyle \frac{8x}{(x^2+4)^2} <0 }\)

OpenStudy (solomonzelman):

Well, we know \(\color{black}{\displaystyle (x^2+4)^2>0 }\) (for all \(\color{black}{\displaystyle x }\)), so in terms of "being grater than zero, all what's going to matter is whether or not \(\color{black}{\displaystyle 8x }\) is negative or not. In other words, whether or not \(\color{black}{\displaystyle 8x<0 }\).

OpenStudy (ny,ny):

8x is less than 0 when x is less than 0(?)

OpenStudy (solomonzelman):

YEs!

OpenStudy (ny,ny):

Ok. So F(x) is decreasing when x<0 ?

OpenStudy (ny,ny):

Or (-infinity,0)

OpenStudy (solomonzelman):

Yes!

OpenStudy (ny,ny):

Thanks